OpenStudy (anonymous):

HELP! solve the equation!!! log3(x+2)+log3(x-4)=3 the 3's after log are little!

OpenStudy (anonymous):

first combine is \[\log_3((x+2)(x-4))=3\] then rewrite in equivalent exponential form as \[(x+2)(x-4)=3^3\] and then it is a quadratic equation to solve don't forget to check your answer once you are done